# Elastic properties and thermodynamic anomalies of supersolids

### v1

@@ I Introduction
Moreover, there has been intensive study on the superfluid fraction and excitation spectrum of the supersolid phase. However, there has been comparatively little study on the elastic properties. In this paper we will build on the homogenisation technique of Rica and co-workers [15, 16, 17] to obtain an effective low-energy theory that agrees well with Bogoliubov theory and numerical calculations. Importantly, the original formulation of the homogenisation technique did not agree with the expected bulk modulus in the superfluid phase, and lacked the internal symmetry constraints for a Cauchy elastic tensor in the supersolid phase. The velocities of the long-wavelength excitations in the previous work also did not agree with the results based on Bogoliubov theory [14].
@@ III Homogenisation for Elasticity
We begin by describing a physical deformation by considering the displacement of points in the material. The displacement can be written as
@@
We then define an inverse passive transformation which maps the coordinate system $\mathbf{r}^{\prime}$ to the coordinate system $\mathbf{r}^{\prime\prime}$, which is a non-Cartesian basis in the material frame that has the same domain as the inital $\mathbf{r}$ (domain $\Omega$). This transformation is of the form
@@
where the basis $\mathbf{r}^{\prime\prime}$ coincides identically with the basis $\mathbf{r}$ in the material frame, i.e. $\mathbf{r}^{\prime\prime}=\mathbf{r}$, but now the space has additional components in its gradients, as well as a Jacobian associated with its volume element. These extra curvatures are necessary to describe the full effect of the displacement procedure on the system. We now relabel the basis $\mathbf{r}^{\prime\prime}$ as $\mathbf{r}$, for convenience, where again we must stress that even though the coordinate systems $\mathbf{r}$ and $\mathbf{r}^{\prime\prime}$ coincide, the space is no longer Cartesian and has an associated Jacobian and curvature. This procedure is illustrated schematically in Figure 2.
@@
As of now, the above transformation is completely general, but in order to proceed we need to specify a particular class of transformation which allows us to use linear response theory and elasticity theory. We make the distinction that $\mathbf{u}({\mathbf{r}})$ is such that all gradients in either space or time are small constants. More specifically, we set $\partial_{i}u_{k}\sim\delta$ and $\partial_{t}u_{k}\sim\delta$, with $\delta$ an arbitrarily small constant that we choose to track up to $\mathop{}\!\mathcal{O}(\delta^{2})$. This is referred to as a uniform deformation, in that the gradient of the deformation is constant everywhere and is small, allowing us to treat it as a perturbative parameter.
@@ IV Elasticity Theory
We now turn to comparing these numerical results from homogenisation with a direct calculation of the elastic constants by finding how the energy of the ground state of the Gross-Pitaevskii equation changes when we deform the simulation cell at fixed particle number (with periodic boundary conditions). We begin with an undeformed ground-state obtained via numerically solving the Gross-Pitaevskii with some given average density, area and interparticle interaction. We then apply a strain to the system by changing the geometry of the simulation cell by a strain tensor $u_{ik}$, whilst keeping the total number of particles fixed, and solving the Gross-Pitaevskii equation for the new geometry to find the deformed ground-state. For sufficiently small strain $u_{ik}$, we expect that such a procedure will yield a ground state with a Lagrangian of the form (20), where the tensor $u_{ik}$ is now a control parameter in our computational experiment.
@@
Our results are presented in Figure 3. These elastic constants should ideally be the same as those obtained via homogenisation theory. In reality, they suffer from numerical errors due to being unable to perform an arbitrarily small strain, as well as convergence errors in the new ground-state when close to the first-order phase transition between the homogeneous superfluid and the supersolid. Scanning across the whole range of the interaction parameter $\Lambda$, we find that, outside of a small region where the phase transition occurs, the elastic constants obtained numerically are in excellent agreement with those obtained analytically, with a typical discrepancy of  0.1%.
In Figure 3, we have also shown the shift in the chemical potential for unit strain: $\mu_{1}^{ik}=\mu_{1}\delta_{ik}$.
@@ V $U(1)$ Phase and Fractional Inertia
In the this section, we consider the application of a phase perturbation to the meanfield supersolid, and analyse the coupling of the phase to the strain field. We follow the metholodology of Josserand et al [15], but our results have some key differences that we will point out.
@@
In order to apply perturbation theory to the phase field, we specify that $\partial_{i}\phi_{0}(\mathbf{r}^{\prime})$ is small and constant within the scope of the unit cell of the supersolid, in a similar way to the strain tensor $u_{ik}$. We note that since the phase is applied in the deformed frame, the total phase in the material frame can be expanded to obtain
@@
which is the final expression for the phase field in the material frame of reference. We can now proceed to expand the Lagrangian in terms of the phase field, using the least action principle to find $\tilde{\phi}$ and the associated energy.
@@
to describe the time derivative of a function which is co-moving with the superfluid velocity field. We can rewrite $\mathbf{A}$ in term of these variables:
@@
Our phase Lagrangian $\mathcal{L}_{\phi}$ in (29) differs from the analogous form in Josserand et al [15] by the existence of the term $\hbar\rho_{0}\partial_{t}\phi_{0}\gradient\cdot\mathbf{u}$. This term is in fact necessary to ensure that $\mathcal{L}_{\phi}$ satisfies Galilean symmetry constraints as pointed out by Son [12]. We also note that this affects the velocity of the excitations of the system, and we will see later in Section VIII in that our results agree with the results obtained in Bogoliubov theory.
Figure 4: The reduction $\varrho/n$ in the superfluid fraction as a function of interaction strength $\Lambda$, exhibiting a discontinuous jump at the transition.
@@ VI Free energy and Grand Canonical ensemble
We assume that the inter-cell number fluctuation is small, and so we can expand the Lagrangian analytically in $\delta n$ up to $\mathop{}\!\mathcal{O}(\delta n^{2})$. We now consider the ground state Lagrangian density $\mathcal{E}(n)=-L_{0}/|\Omega|$ which is a function of the average density $n$ of the undeformed system. Under a change of the local density, this can be expanded as
@@ VII Effective supersolid Lagrangian
It is interesting to note that, although we expect the Gross-Pitaevskii Lagrangian to struggle with describing a normal solid, the effective phenomenological Lagrangian that emerges actually returns to that of a normal solid in the limit $\varrho\rightarrow n$.
@@
Figure 5: Top: excitation velocities for all available sound modes of a system governed by a Heaviside interaction of strength $\Lambda$ with a phase transition at $\Lambda\sim 38$, as calculated via homogenisation. Bottom: the eigenvector angle $\theta$ for the available longitudinal modes.
@@ IX Results and Discussion
$\mathcal{R}=\frac{B_{e}}{B_{t}}\,.$ (63)
@@
Figure 7: Compressibility ratio $\mathcal{R}$ for a system governed by a soft-core interaction of strength $\Lambda$. The expected value of $\mathcal{R}=1$ is obtained in the superfluid phase, but is starkly differenti n the supersolid regime.
@@ Appendix D Numerical Deformations
$\mathbf{V}=\begin{pmatrix}\epsilon&0\\ 0&0\end{pmatrix}\!:F_{2}$ $=P\epsilon+\left(A_{xxxx}+\frac{A_{xyxy}}{3}\right)\frac{\epsilon^{2}}{2}$ (110)
@@
$\mathbf{V}=\begin{pmatrix}\epsilon&0\\ 0&\epsilon\end{pmatrix}\!:F_{2}$ $=2P\epsilon+\left(\frac{A_{xxxx}+A_{yyyy}}{2}+A_{xxyy}\right)\epsilon^{2}\,.$ (116)

### v2

@@ I Introduction
Moreover, there has been intensive study on the superfluid fraction and excitation spectrum of the supersolid phase. However, there has been comparatively little study on the elastic properties. In this paper we will build on the homogenisation technique of Rica and co-workers [15, 16, 17] to obtain an effective low-energy theory that agrees well with Bogoliubov theory and numerical calculations. Importantly, the original formulation of the homogenisation technique obtained a value for the bulk modulus that was not in agreement with the expected bulk modulus in the superfluid phase. Furthermore, a triangular supersolid would be expected to possess certain symmetries in the Cauchy elastic tensor, and these were not satisfied in the original formulation. Finally, the velocities of the long-wavelength excitations in the previous work also did not agree with the results based on Bogoliubov theory [14], in the sense that the velocities obtained on a particular supersolid phase using the Bogoliubov technique were not consistent with those obtained via the original formulation of homogenisation.
@@ III Homogenisation for Elasticity
We begin by describing a physical deformation by considering the displacement, $u(r)$, of points in the material. The displacement can be written as
@@
We then define an inverse passive transformation which maps the coordinate system $\mathbf{r}^{\prime}$ to the coordinate system $\mathbf{r}^{\prime\prime}$, which is a non-Cartesian basis in the material frame that has the same domain as the inital $\mathbf{r}$ (domain $\Omega$). This passive transformation is defined by the displacement $\mathbf{u}^{\prime}(\mathbf{r}^{\prime})$, where we note that this does not change the energy of the system in any way. It is simply a redefinition of coordinates in order for us to express the actual change in the energy that occured during the active transformation defined in (5).
This passive transformation is given by the form
@@
where the basis $\mathbf{r}^{\prime\prime}$ coincides identically with the basis $\mathbf{r}$ in the material frame, i.e. $\mathbf{r}^{\prime\prime}=\mathbf{r}$, but now the space has additional components in its gradients, as well as a Jacobian associated with its volume element. Note that the passive transformation $\mathbf{u}^{\prime}(\mathbf{r}^{\prime})$ is in the deformed lab frame given by the coordinates $\mathbf{r}^{\prime}$, and importantly is a different transformation to $\mathbf{u}(\mathbf{r})$.
These extra curvatures are necessary to describe the full effect of the displacement procedure on the system. We now relabel the basis $\mathbf{r}^{\prime\prime}$ as $\mathbf{r}$, for convenience, where again we must stress that even though the coordinate systems $\mathbf{r}$ and $\mathbf{r}^{\prime\prime}$ coincide, the space is no longer Cartesian and has an associated Jacobian and curvature. This procedure is illustrated schematically in Figure 2.
@@
As of now, the above transformation is completely general, but in order to proceed we need to specify a particular class of transformation which allows us to use linear response theory and elasticity theory. We make the distinction that $\mathbf{u}({\mathbf{r}})$ is such that all gradients in space are small constants and that gradients in time are considered small velocities.
More specifically, we choose the tensor $\partial_{i}u_{k}$, which is a dimensionless number, s.t. $\partial_{i}u_{k}\sim\delta$, where $\delta\ll 1$. The time derivative of $\mathbf{u}$, i.e. $\partial_{t}u_{k}$, is treated as small in the same sense as the phase perturbation, which we will later cover in Section V. This is referred to as a uniform deformation, in that the strain tensor is constant and small everywhere, which allows us to treat it as a perturbative parameter which we will expand up to harmonic terms.
@@ IV Elasticity Theory
We now turn to comparing these numerical results from homogenisation with a direct calculation of the elastic constants by finding how the energy of the ground state of the Gross-Pitaevskii equation changes when we deform the simulation cell at fixed particle number (with periodic boundary conditions). We begin with an undeformed ground-state obtained via numerically solving the Gross-Pitaevskii with some given average density, area and interparticle interaction. We then apply a small strain to the system by changing the geometry of the simulation cell by a strain tensor $u_{ik}$, whilst keeping the total number of particles fixed, and solving the Gross-Pitaevskii equation for the new geometry to find the deformed ground-state. For sufficiently small strain $u_{ik}$, we expect that such a procedure will yield a ground state with a Lagrangian of the form (20), where the tensor $u_{ik}$ is now a control parameter in our computational experiment. This strain is small in the same sense as defined in the expansion of the Lagrangian. More precisely, we consider $u_{ik}\ll 1$, and then try to find the harmonic terms in the energy perturbation.
@@
Our results are presented in Figure 3, where we have plotted the bulk and shear modulus $K$ and $G$, the pressure $P$ and the shift in the chemical potential for unit strain: $\mu_{1}^{ik}=\mu_{1}\delta_{ik}$. These elastic constants should ideally be the same as those obtained via homogenisation theory. In reality, they suffer from numerical errors due to being unable to perform an arbitrarily small strain, as well as convergence errors in the new ground-state when close to the first-order phase transition between the homogeneous superfluid and the supersolid. Scanning across the whole range of the interaction parameter $\Lambda$, we find that, outside of a small region where the phase transition occurs, the elastic constants obtained numerically are in excellent agreement with those obtained analytically, with a typical discrepancy of $\sim 0.1\%$.
We found that an applied numerical strain which was smaller than $\sim 5\%$ (by this we mean $u_{ik}\sim 0.05$) would generate this excellent agreement, but if the applied strain was $\sim 10\%$, the numerical value would diverge from that of homogenisation. We suspect that this is due to the introduction of anharmonic elastic moduli which are now no longer negligible.
@@ V $U(1)$ Phase and Fractional Inertia
In this section, we consider the application of a phase perturbation to the meanfield supersolid, and analyse the coupling of the phase to the strain field. We follow the metholodology of Josserand et al [15], but our results have some key differences that we will point out.
@@
In order to apply perturbation theory to the phase field, we specify that $\partial_{i}\phi_{0}(\mathbf{r}^{\prime})$ is small and constant within the scope of the unit cell of the supersolid, in a similar way to the strain tensor $u_{ik}$. This is equivalent to stating that the phase change across the unit cell is much smaller than $2\pi$, i.e. $\Delta\phi\ll 2\pi$. Note also that we are not considering the generation of vortices.
We note that since the phase is applied in the deformed frame, the total phase in the material frame can be expanded to obtain
@@
where $\mathbf{u}$ is the displacement as defined in (5). This is the final expression for the phase field in the material frame of reference. We can now proceed to expand the Lagrangian in terms of the phase field, using the least action principle to find $\tilde{\phi}$ and the associated energy.
@@
to describe the time derivative of a function which is co-moving with the superfluid velocity field.
Note here the appearance of the term $m/\hbar\partial_{t}\mathbf{u}$ which we previously stated as ‘small’ without further elaboration. Here we define this term to be small on the same order as $\mathbf{\nabla}\phi_{0}$, i.e. that the velocity induced by the elastic deformation is comparable to the velocity induced by the phase change.
We can rewrite $\mathbf{A}$ in term of these variables:
@@
Our phase Lagrangian $\mathcal{L}_{\phi}$ in (29) differs from the analogous form in Josserand et al [15] by the existence of the term $\hbar\rho_{0}\partial_{t}\phi_{0}\gradient\cdot\mathbf{u}$. This term is in fact necessary to ensure that $\mathcal{L}_{\phi}$ satisfies Galilean symmetry constraints as pointed out by Son [12]. We also note that this affects the velocity of the excitations of the system, specifically in reference to (46) where this term is directly responsible for the $n$ component o $\hbar\left(n-\varrho+\mu_{1}/\mathcal{E}^{\prime\prime}\right)\dot{\phi}\mathbf{\nabla}\cdot\mathbf{u}$. If one we re to calculate th velocities without this critical term, one would see that it differs significantly from those predicted by Bogoliubov theory (shown in Section VIII).
Figure 4: The superfluid fraction $f_{s}\equiv 1-\varrho/n$ as a function of $\Lambda$. We appear to observe a discontinuous transition at $\Lambda\sim 38$, which remains even after checking the transition region with a much finer $\Lambda$ discretisation. This appears to be numerical verification of [17], which predicted a first order transition close to this value.
@@ VI Free energy and Grand Canonical ensemble
We assume that the inter-cell number fluctuation is small, i.e. that $\delta n/n\ll 1$, and so we can expand the Lagrangian analytically in $\delta n$ up to $\mathop{}\!\mathcal{O}(\delta n^{2})$. We now consider the ground state Lagrangian density $\mathcal{E}(n)=-L_{0}/|\Omega|$ which is a function of the average density $n$ of the undeformed system. Under a change of the local density, this can be expanded as
@@ VII Effective supersolid Lagrangian
Note that for the following derivation of the supersolid Goldstone velocities, it is assumed that the ground state is a supersolid and hence $\varrho\neq 0$. In a superfluid state ($\varrho=0$), the following analysis does not apply, and instead one needs to refer to (47).
@@
It is interesting to note that, although we expect the Gross-Pitaevskii Lagrangian to struggle with describing a normal solid, the effective phenomenological Lagrangian that emerges actually returns to that of a normal solid in the limit $\varrho\rightarrow n$. Moreover, although all of the modes have a factor of $1/\varrho$ in their velocities which seems to indicate that the velocities should blow up as $\varrho\rightarrow 0$, the velocities appear to be well behaved as $\varrho$ becomes small. This suggests that there is some implicit dependence on $\varrho$ hidden in the values of $B_{iklm}$ etc., which stops the velocities from blowing up. However, we should note that as we appear to see a first-order transition, the value of $\varrho$ never really approaches zero in our numerics, but only jumps directly to zero at which point these equations are no longer valid.
@@
Figure 5: Top: excitation velocities (in units of $\hbar/\sqrt{n}ma^{2}$) for all available sound modes of a system governed by a Heaviside interaction of strength $\Lambda$ with a phase transition at $\Lambda\sim 38$, as calculated via homogenisation. Note that in the superfluid phase the bulk mode is the Bogoliubov mode, hence the sudden appearance of two extra modes at the superfluid-supersolid transtion. Bottom: the eigenvector angle $\theta$ for the available longitudinal modes.
@@ IX Results and Discussion
$\mathcal{R}=\frac{B_{e}}{B_{t}}\equiv\frac{K}{n\mu_{1}}\,.$ (63)
@@
Figure 7: Compressibility ratio $\mathcal{R}$ for a system governed by a soft-core interaction of strength $\Lambda$. $\mathcal{R}$ is obtained using the relation (63), where the values are calculated using homogenisation. The expected value of $\mathcal{R}=1$ is obtained in the superfluid phase, but is starkly different in the supersolid regime.
@@ Appendix D Numerical Deformations
$\mathbf{V}=\begin{pmatrix}\epsilon&0\\ 0&0\end{pmatrix}\!:F_{2}$ $=P\epsilon+\left(A_{xxxx}+\frac{A_{xyxy}}{3}\right)\frac{\epsilon^{2}}{2}$
@@
$\mathbf{V}=\begin{pmatrix}\epsilon&0\\ 0&\epsilon\end{pmatrix}\!:F_{2}$ $=2P\epsilon+\left(\frac{A_{xxxx}+A_{yyyy}}{2}+A_{xxyy}\right)\epsilon^{2}\,.$